OpenStudy (lucyluv16):

Simplify square root parenthesis 1 minus sine theta parenthesis times parenthesis 1 plus sine theta parenthesis.

OpenStudy (lucyluv16):

@MrNood @phi please help!!!!

OpenStudy (lucyluv16):

\[\sqrt{(1-\sin \theta)(1+\sin \theta)}\]

OpenStudy (lucyluv16):

idk how to solve it!!!

OpenStudy (lucyluv16):

A - ±sinθ B - cos θ C- ±tan θ D- \[\sqrt{\sin}\]

OpenStudy (lucyluv16):

theta, sqrt sin theta for D

OpenStudy (lucyluv16):

@princesskitty01

OpenStudy (phi):

do you know the "pattern" (a-b)(a+b) = a^2 - b^2 (You can use FOIL to show this is true) anyway, if you use that, what does the inside of the square root become ?

OpenStudy (lucyluv16):

I don't know that doesn't make sense to me..... @phi

OpenStudy (lucyluv16):

@DivineSolar

OpenStudy (phi):

it should. You should have learned how to factor a "difference of squares" i.e. x^2 -4 factors into (x-2)(x+2) you should also have learned how to multiply (x-y)(x+y) to get x^2 - y^2 or if all that is mysterious you should know how to distribute for example A(1+ sin) = A + A sin now if A is in fact (1-sin) we get 1-sin+ (1-sin)*sin now distribute the 2nd part: 1-sin + sin - sin*sin or 1 - sin^2

OpenStudy (lucyluv16):

ok that I understand, so would the answer be- \[\sqrt{\sin \theta}\] ???? @phi

OpenStudy (lucyluv16):

wait no it wouldn't because the theta would have been cancled out, would it actually be +- sin????

OpenStudy (phi):

how did you get just sin ? (btw, I'm leaving out the theta because it's too hard to type) re-read what I posted. the inside is not just sin

OpenStudy (phi):

you should get \[ \sqrt{1 - \sin^2 \theta }\]

OpenStudy (phi):

and that can be simplified using sin^2 + cos^2 = 1 which means cos^2 = 1 - sin^2

OpenStudy (lucyluv16):

ohhhhhhhhh, ok, I understand that

OpenStudy (lucyluv16):

can you help me with one other problem? I just want to make sure I got it correct or not

OpenStudy (phi):

first, what did you get for an answer to this question ?

OpenStudy (lucyluv16):

cos

OpenStudy (lucyluv16):

right.....?......

OpenStudy (phi):

yes

OpenStudy (lucyluv16):

awesome can u help me with one more/?? @phi